Let τ be a fixed odd permutation in . S₁₀ Show that every odd permutation in S₁₀ is

a product of τ and some permutation in A₁₀


1
Expert's answer
2021-06-07T13:38:03-0400

Solution:

Proof. Let "\sigma^{\prime}" be an odd permutation in "S_{n}" . We must show that there exists an even permutation "\mu \in A_{n}" such that "\sigma^{\prime}=\sigma \mu" . Indeed, we may take "\mu=\sigma^{-1} \sigma^{\prime}" , since as the product of two odd permutations, it is an even permutation, and

"\sigma^{\prime}=\sigma\left(\sigma^{-1} \sigma^{\prime}\right)"

For completeness, let's prove directly that "\sigma^{-1} \sigma^{\prime}" is even. From the definition of an odd permutation, there exist a finite number of transpositions "\tau_{1}, \ldots, \tau_{m}" for some odd "m \in \mathbb{N}" such that

"\sigma=\tau_{1} \ldots \tau_{m}"

Similarly, since "\sigma^{\prime}" is also an odd permutation, there exist a finite number of transpositions "\tau_{1}^{\prime}, \ldots, \tau_{\ell}^{\prime}" for some odd "\ell \in \mathbb{N}" such that "\sigma^{\prime}=\tau_{1}^{\prime} \ldots \tau_{\ell}^{\prime}" . Consider now the permutation

"\mu=\sigma^{-1} \sigma^{\prime}"

I claim that this lies in "A_{n}" . Indeed we have

"\mu=\sigma^{-1} \sigma^{\prime}=\underbrace{\tau_{m} \ldots \tau_{1} \tau_{1}^{\prime} \ldots \tau_{\ell}^{\prime}}_{m+\ell}"

The sum of two odd numbers is even, and so it follows that this is an even permutation.

Thus, if we take "n=10" , we can prove it same way.